# Gravitational wave propagation in bigravity in the late universe

### v1

@@ Acknowledgments
We are grateful to S. Castrignano, R. Oliveri, and A. Tagliacozzo for helpful discussions. ASO acknowledges financial support from the fellowship PIF21/237 of the UPV/EHU. MdC acknowledges support from INFN iniziativa specifica GeoSymQFT. This work has been supported by the Basque Government Grant IT1628-22 and by the Grant PID2021-123226NB-I00 (funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”).
@@ Appendix A Asymptotic expansion of the $v$ mode
For completeness, it is also worth mentioning that, if desired, one can obtain an approximation of the exact solution (3.7b) for $\xi\geq 0$ (real and non-negative) in the sub-horizon regime, characterized by $k|\eta|\gg 1$, which corresponds to the observable modes in the present universe. This case is going to be restricted to a very narrow set of theoretically allowed mass values due to the Higuchi bound $m_{\rm FP}^{2}>2H^{2}$, corresponding to admissible values in the interval $20\leq\xi<1/$ 2. In this regime, one can use the asymptotic form of Bessel functions for fixed real and non-negative order and large arguments [45],
@@ B.1 Autocorrelation of $u$
$\begin{split}&\langle u(\eta,x)u(\eta^{\prime},x^{\prime})\rangle-\langle u(\eta,x)\rangle\langle u(\eta^{\prime},x^{\prime})\rangle=\\ &\left(\frac{2\cos\theta}{M_{g}^{2}}\right)^{2}\int\limits_{\eta_{0}}^{\eta}\mbox{d}\bar{\eta}\int\limits_{\eta_{0}}^{\eta^{\prime}}\mbox{d}\bar{\bar{\eta}}\int\limits_{-\infty}^{+\infty}\frac{\mbox{d}k_{x}}{2\pi}\int\limits_{-\infty}^{+\infty}\frac{\mbox{d}k^{\prime}_{x}}{2\pi}\,\frac{\sin(k_{x}(\eta-\bar{\eta}))}{k_{x}}\frac{\sin(k^{\prime}_{x}(\eta^{\prime}-\bar{\bar{\eta}}))}{k^{\prime}_{x}}e^{i(k_{x}x+k^{\prime}_{x}x^{\prime})}\langle\pi(k,\bar{\eta})\pi(k^{\prime},\bar{\bar{\eta}})\rangle=\\ &\left(\frac{2s\cos\theta}{M_{g}^{2}}\right)^{2}\int\limits_{\eta_{0}}^{\eta}\mbox{d}\bar{\eta}\int\limits_{\eta_{0}}^{\eta^{\prime}}\mbox{d}\bar{\bar{\eta}}\int\limits_{-\infty}^{+\infty}\frac{\mbox{d}k_{x}}{2\pi}\,\frac{\sin(k_{x}(\eta-\bar{\eta}))}{k_{x}}\frac{\sin(k_{x}(\eta^{\prime}-\bar{\bar{\eta}}))}{k_{x}}e^{ik_{x}(x-x^{\prime})}\delta(\bar{\eta}-\bar{\bar{\eta}})=\\ &\left(\frac{2s\cos\theta}{M_{g}^{2}}\right)^{2}\Theta(\eta-\eta^{\prime})\int\limits_{\eta_{0}}^{\eta^{\prime}}\mbox{d}\bar{\eta}\int\limits_{-\infty}^{+\infty}\frac{\mbox{d}k_{x}}{2\pi}\,\frac{\sin(k_{x}(\eta-\bar{\eta}))\sin(k_{x}(\eta^{\prime}-\bar{\eta}))}{k_{x}^{2}}e^{ik_{x}(x-x^{\prime})}+(\eta\stackrel{{\scriptstyle\text{sym}}}{{\longleftrightarrow}}\eta^{\prime})=\\ &-\frac{1}{8}\left(\frac{s\cos\theta}{M_{g}^{2}}\right)^{2}\Theta(\eta-\eta^{\prime})\Big{[}4(\eta^{\prime}-\eta_{0})\big{(}\left|x-x^{\prime}+\eta-\eta^{\prime}\right|+\left|x-x^{\prime}-\eta+\eta^{\prime}\right|\big{)}+\\ &(\eta-\eta^{\prime}+x-x^{\prime})|\eta-\eta^{\prime}+x-x^{\prime}|-(\eta^{\prime}-\eta+x-x^{\prime})|\eta^{\prime}-\eta+x-x^{\prime}|-\\ &(\eta-2\eta_{0}+\eta^{\prime}+x-x^{\prime})|\eta-2\eta_{0}+\eta^{\prime}+x-x^{\prime}|-(\eta-2\eta_{0}+\eta^{\prime}+x^{\prime}-x)|\eta-2\eta_{0}+\eta^{\prime}+x^{\prime}-x|\Big{]}+(\eta\stackrel{{\scriptstyle\text{sym}}}{{\longleftrightarrow}}\eta^{\prime})~{}.\end{split}$ (B.1)
@@
$\begin{split}&\text{R}_{uu}\big{(}\eta,x;\eta^{\prime},x^{\prime}\big{)}=\frac{\langle u(\eta,x)u(\eta^{\prime},x^{\prime})\rangle-\langle u(\eta,x)\rangle\langle u(\eta^{\prime},x^{\prime})\rangle}{\sqrt{{\rm Var}\big{[}u(\eta,x)\big{]}{\rm Var}\big{[}u(\eta^{\prime},x^{\prime})\big{]}}}=\\ &-\frac{1}{8(\eta-\eta_{0})(\eta^{\prime}-\eta_{0})}\Big{[}4(\eta^{\prime}-\eta_{0})\big{(}\left|x-x^{\prime}+\eta-\eta^{\prime}\right|+\left|x-x^{\prime}-\eta+\eta^{\prime}\right|\big{)}+(\eta-\eta^{\prime}+x-x^{\prime})|\eta-\eta^{\prime}+x-x^{\prime}|\\ &-(\eta^{\prime}-\eta+x-x^{\prime})|\eta^{\prime}-\eta+x-x^{\prime}|-(\eta-2\eta_{0}+\eta^{\prime}+x-x^{\prime})|\eta-2\eta_{0}+\eta^{\prime}+x-x^{\prime}|\\ &-(\eta-2\eta_{0}+\eta^{\prime}+x^{\prime}-x)|\eta-2\eta_{0}+\eta^{\prime}+x^{\prime}-x|\Big{]}\Theta(\eta-\eta^{\prime})+(\eta\stackrel{{\scriptstyle\text{sym}}}{{\longleftrightarrow}}\eta^{\prime})~{}.\end{split}$ (B.3)
@@ B.2 Autocorrelation of $v$
$\begin{split}&\langle v(\eta,x)v(\eta^{\prime},x^{\prime})\rangle-\langle v(\eta,x)\rangle\langle v(\eta^{\prime},x^{\prime})\rangle=\\ &\left(\frac{2s\sin\theta}{M_{g}^{2}}\right)^{2}\int\limits_{\eta_{0}}^{\eta}\mbox{d}\bar{\eta}\int\limits_{\eta_{0}}^{\eta^{\prime}}\mbox{d}\bar{\bar{\eta}}\int\limits_{-\infty}^{+\infty}\frac{\mbox{d}k_{x}}{2\pi}\,\frac{\sin(\omega_{k}(\eta-\bar{\eta}))}{\omega_{k}}\frac{\sin(\omega_{k}(\eta^{\prime}-\bar{\bar{\eta}}))}{\omega_{k}}e^{ik_{x}(x-x^{\prime})}\delta(\bar{\eta}-\bar{\bar{\eta}})=\\ &\left(\frac{2s\sin\theta}{M_{g}^{2}}\right)^{2}\Theta(\eta-\eta^{\prime})\int\limits_{\eta_{0}}^{\eta^{\prime}}\mbox{d}\bar{\eta}\int\limits_{-\infty}^{+\infty}\frac{\mbox{d}k_{x}}{2\pi}\,\frac{\sin(\omega_{k}(\eta-\bar{\eta}))\sin(\omega_{k}(\eta^{\prime}-\bar{\eta}))}{\omega_{k}^{2}}e^{ik_{x}(x-x^{\prime})}+(\eta\stackrel{{\scriptstyle\text{sym}}}{{\longleftrightarrow}}\eta^{\prime})=\\ &\left(\frac{2s\sin\theta}{M_{g}^{2}}\right)^{2}\Theta(\eta-\eta^{\prime})\int\limits_{-\infty}^{+\infty}\text{d}k_{x}\;e^{ik_{x}(x-x^{\prime})}\Bigg{(}\frac{(\eta^{\prime}-\eta_{0})\cos{\left(\sqrt{k_{x}^{2}+m_{\rm FP}^{2}}(\eta-\eta^{\prime})\right)}}{4\pi(k_{x}^{2}+m_{\rm FP}^{2})}\\ &+\frac{\sin{\left(\sqrt{k_{x}^{2}+m_{\rm FP}^{2}}(\eta-\eta^{\prime})\right)}-\sin{\left(\sqrt{k_{x}^{2}+m_{\rm FP}^{2}}(\eta+\eta^{\prime}-2\eta_{0})\right)}}{8\pi(k_{x}^{2}+m_{\rm FP}^{2})^{3/2}}\Bigg{)}+(\eta\stackrel{{\scriptstyle\text{sym}}}{{\longleftrightarrow}}\eta^{\prime})~{}.\end{split}$ (B.4)
@@
$\begin{split}&\int\limits_{-m_{\rm FP}}^{m_{\rm FP}}\text{d}u\,\dfrac{e^{-u\,q}}{m_{\rm FP}^{2}-u^{2}}\sin{\left(\sqrt{m_{\rm FP}^{2}-u^{2}}|p|\right)}\approx\\ &\quad\frac{1}{m_{\rm FP}}\Bigg{[}\pi\cosh(m_{\rm FP}|q|)-\sqrt{\frac{2\pi}{m_{\rm FP}|p|}}\left(1-\frac{q^{2}}{p^{2}}\right)^{1/4}\cos\Big{(}m_{\rm FP}\sqrt{p^{2}-q^{2}}+\frac{\pi}{4}\Big{)}\Bigg{]}~{},\;\;\mbox{for $|p|>|q|$\,,\; $m_{\rm FP}|p|\gg 1$}~{}.\end{split}$ (B.11)
@@
    $\begin{split}&\langle v(\eta,x)v(\eta^{\prime},x^{\prime})\rangle-\langle v(\eta,x)\rangle\langle v(\eta^{\prime},x^{\prime})\rangle\\ &\approx\left(\frac{2s\sin\theta}{M_{g}^{2}}\right)^{2}\frac{(\eta^{\prime}-\eta_{0})\left((\eta-\eta^{\prime})^{2}-\Delta x^{2}\right)^{1/4}\cos{\left(\frac{\pi}{4}+m_{\rm FP}\sqrt{(\eta-\eta^{\prime})^{2}-\Delta x^{2}}\right)}}{2\sqrt{2\pi\,m_{\rm FP}^{3}}(\eta-\eta^{\prime})}\Theta(\eta-\eta^{\prime})\\ &\quad+\,(\eta\stackrel{{\scriptstyle\text{sym}}}{{\longleftrightarrow}}\eta^{\prime})\end{split}$ (B.15)
@@
    $\begin{split}&\langle v(\eta,x)v(\eta^{\prime},x^{\prime})\rangle-\langle v(\eta,x)\rangle\langle v(\eta^{\prime},x^{\prime})\rangle\\ &\approx-\left(\frac{2s\sin\theta}{M_{g}^{2}}\right)^{2}\frac{\left((\eta+\eta^{\prime}-2\eta_{0})^{2}-\Delta x^{2}\right)^{3/4}\sin{\left(\frac{\pi}{4}+m_{\rm FP}\sqrt{(\eta+\eta^{\prime}-2\eta_{0})^{2}-\Delta x^{2}}\right)}}{4\sqrt{2\pi m_{\rm FP}^{5}}(\eta+\eta^{\prime}-2\eta_{0})^{2}}\Theta(\eta-\eta^{\prime})\\ &\quad+\,(\eta\stackrel{{\scriptstyle\text{sym}}}{{\longleftrightarrow}}\eta^{\prime})~{}.\end{split}$ (B.16)
@@ B.3 Correlation between $u$ and $v$
$\begin{split}&\langle u(\eta,x)v(\eta^{\prime},x^{\prime})\rangle-\langle u(\eta,x)\rangle\langle v(\eta^{\prime},x^{\prime})\rangle=\\ &-\frac{2s^{2}\sin(2\theta)}{M_{g}^{4}}\Theta(\eta-\eta^{\prime})\int\limits_{\eta_{0}}^{\eta^{\prime}}\mbox{d}\bar{\eta}\int\limits_{-\infty}^{+\infty}\frac{\mbox{d}k_{x}}{2\pi}\,\frac{\sin(k_{x}(\eta-\bar{\eta}))\sin(\omega_{k}(\eta^{\prime}-\bar{\eta}))}{k_{x}\,\omega_{k}}e^{ik_{x}(x-x^{\prime})}+(\eta\stackrel{{\scriptstyle\text{sym}}}{{\longleftrightarrow}}\eta^{\prime})\\ &=-\frac{2s^{2}\sin(2\theta)}{M_{g}^{4}}\Theta(\eta-\eta^{\prime})\int\limits_{-\infty}^{+\infty}\frac{\mbox{d}k_{x}}{2\pi m_{\rm FP}^{2}}\Bigg{(}\frac{\sin{\left(k_{x}(\eta-\eta^{\prime})\right)-\cos{\left(\sqrt{k_{x}^{2}+m_{\rm FP}^{2}}(\eta^{\prime}-\eta_{0})\right)}\sin{\left(k_{x}(\eta-\eta_{0})\right)}}}{k}\\ &+\frac{\cos{\left(k_{x}(\eta-\eta_{0})\right)}\sin{\left(\sqrt{k_{x}^{2}+m_{\rm FP}^{2}}(\eta^{\prime}-\eta_{0})\right)}}{\sqrt{k_{x}^{2}+m_{\rm FP}^{2}}}\Bigg{)}e^{ik_{x}(x-x^{\prime})}+(\eta\stackrel{{\scriptstyle\text{sym}}}{{\longleftrightarrow}}\eta^{\prime})~{},\end{split}$ (B.19)
@@
$\begin{split}&\int\limits_{-\infty}^{+\infty}\frac{\mbox{d}k_{x}}{2\pi m_{\rm FP}^{2}}\Bigg{(}\frac{\sin{\left(k_{x}(\eta-\eta^{\prime})\right)-\cos{\left(\sqrt{k_{x}^{2}+m_{\rm FP}^{2}}(\eta^{\prime}-\eta_{0})\right)}\sin{\left(k_{x}(\eta-\eta_{0})\right)}}}{k_{x}}\\ &\quad+\frac{\cos{\left(k(\eta-\eta_{0})\right)}\sin{\left(\sqrt{k_{x}^{2}+m_{\rm FP}^{2}}(\eta^{\prime}-\eta_{0})\right)}}{\sqrt{k_{x}^{2}+m_{\rm FP}^{2}}}\Bigg{)}e^{ik_{x}(x-x^{\prime})}\\ &=\frac{\text{sgn}(x-x^{\prime}+\eta-\eta^{\prime})-\text{sgn}(x-x^{\prime}-\eta+\eta^{\prime})}{4m_{\rm FP}^{2}}+\frac{i}{8\pi\,m_{\rm FP}^{2}}\Big{[}J(\Delta\eta^{\prime},\nu)+J(-\Delta\eta^{\prime},\nu)\\ &\quad-J(\Delta\eta^{\prime},\tilde{\nu})-J(-\Delta\eta^{\prime},\tilde{\nu})-I_{1}(\Delta\eta^{\prime},\nu)+I_{1}(-\Delta\eta^{\prime},\nu)-I_{1}(\Delta\eta^{\prime},\tilde{\nu})+I_{1}(-\Delta\eta^{\prime},\tilde{\nu})\Big{]}~{},\end{split}$ (B.20)
@@
    $\begin{split}&\langle u(\eta,x)v(\eta^{\prime},x^{\prime})\rangle-\langle u(\eta,x)\rangle\langle v(\eta^{\prime},x^{\prime})\rangle\\ &=-\frac{2s^{2}\sin(2\theta)}{M_{g}^{4}}\frac{\text{sgn}(\Delta x+(\eta-\eta^{\prime}))-\text{sgn}(\Delta x-(\eta-\eta^{\prime}))-2\cos{\left(m_{\rm FP}(\eta^{\prime}-\eta_{0})\right)}}{4m_{\rm FP}^{2}}\Theta(\eta-\eta^{\prime})\\ &\quad+\,(\eta\stackrel{{\scriptstyle\text{sym}}}{{\longleftrightarrow}}\eta^{\prime})~{}.\end{split}$ (B.24)
@@
    $\begin{split}&\langle u(\eta,x)v(\eta^{\prime},x^{\prime})\rangle-\langle u(\eta,x)\rangle\langle v(\eta^{\prime},x^{\prime})\rangle\\ &\approx-\frac{2s^{2}\sin(2\theta)}{M_{g}^{4}}\frac{(\eta+\eta^{\prime}-2\eta_{0}-|\Delta x|)\sin{\left(\frac{\pi}{4}+m_{\rm FP}\sqrt{(\eta^{\prime}-\eta_{0})^{2}-(\eta-\eta_{0}-|\Delta x|)^{2}}\right)}}{2\sqrt{2\pi\,m_{\rm FP}^{5}}(\eta-\eta_{0}-|\Delta x|)[(|\Delta x|-(\eta-\eta^{\prime}))(\eta+\eta^{\prime}-2\eta_{0}-|\Delta x|)]^{1/4}}\Theta(\eta-\eta^{\prime})\\ &\quad+\,(\eta\stackrel{{\scriptstyle\text{sym}}}{{\longleftrightarrow}}\eta^{\prime})~{}.\end{split}$ (B.25)
@@
  • spacelike separated events without a common causal past,$|x-x^{\prime}|\geq\eta+\eta^{\prime}-2\eta_{0}>\eta-\eta^{\prime}$:
@@
[…]

### v2

@@ Acknowledgments
We are grateful to S. Castrignano, R. Oliveri, I.D. Saltas, and A. Tagliacozzo for helpful discussions, and to an anonymous referee for their constructive feedback. ASO acknowledges financial support from the fellowship PIF21/237 of the UPV/EHU. MdC acknowledges support from INFN iniziativa specifica GeoSymQFT. This work has been supported by the Basque Government Grant IT1628-22 and by the Grant PID2021-123226NB-I00 (funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”).
@@ Appendix A Asymptotic expansion of the $v$ mode
For completeness, it is also worth mentioning that, if desired, one can obtain an approximation of the exact solution (3.7b) for $\xi\geq 0$ (real and non-negative) in the sub-horizon regime, characterized by $k|\eta|\gg 1$, which corresponds to the observable modes in the present universe. This case is going to be restricted to a very narrow set of theoretically allowed mass values due to the Higuchi bound $m_{\rm FP}^{2}>2H^{2}$, corresponding to admissible values in the interval $0\leq\xi<1/2$. In this regime, one can use the asymptotic form of Bessel functions for fixed real and non-negative order and large arguments [62],
@@ B.1 Autocorrelation of $u$
$\begin{split}&\langle u(\eta,x)u(\eta^{\prime},x^{\prime})\rangle-\langle u(\eta,x)\rangle\langle u(\eta^{\prime},x^{\prime})\rangle=\\ &\left(\frac{2\cos\theta}{M_{g}^{2}}\right)^{2}\int\limits_{\eta_{0}}^{\eta}\mbox{d}\bar{\eta}\int\limits_{\eta_{0}}^{\eta^{\prime}}\mbox{d}\bar{\bar{\eta}}\int\limits_{-\infty}^{+\infty}\frac{\mbox{d}k_{x}}{2\pi}\int\limits_{-\infty}^{+\infty}\frac{\mbox{d}k^{\prime}_{x}}{2\pi}\,\frac{\sin(k_{x}(\eta-\bar{\eta}))}{k_{x}}\frac{\sin(k^{\prime}_{x}(\eta^{\prime}-\bar{\bar{\eta}}))}{k^{\prime}_{x}}e^{i(k_{x}x+k^{\prime}_{x}x^{\prime})}\langle\pi(k,\bar{\eta})\pi(k^{\prime},\bar{\bar{\eta}})\rangle=\\ &\left(\frac{2s\cos\theta}{M_{g}^{2}}\right)^{2}\int\limits_{\eta_{0}}^{\eta}\mbox{d}\bar{\eta}\int\limits_{\eta_{0}}^{\eta^{\prime}}\mbox{d}\bar{\bar{\eta}}\int\limits_{-\infty}^{+\infty}\frac{\mbox{d}k_{x}}{2\pi}\,\frac{\sin(k_{x}(\eta-\bar{\eta}))}{k_{x}}\frac{\sin(k_{x}(\eta^{\prime}-\bar{\bar{\eta}}))}{k_{x}}e^{ik_{x}(x-x^{\prime})}\delta(\bar{\eta}-\bar{\bar{\eta}})=\\ &\left(\frac{2s\cos\theta}{M_{g}^{2}}\right)^{2}\Theta(\eta-\eta^{\prime})\int\limits_{\eta_{0}}^{\eta^{\prime}}\mbox{d}\bar{\eta}\int\limits_{-\infty}^{+\infty}\frac{\mbox{d}k_{x}}{2\pi}\,\frac{\sin(k_{x}(\eta-\bar{\eta}))\sin(k_{x}(\eta^{\prime}-\bar{\eta}))}{k_{x}^{2}}e^{ik_{x}(x-x^{\prime})}+(\eta\mathrel{{\mathop{\longleftrightarrow}\limits^{\text{sym}}}}\eta^{\prime})=\\ &-\frac{1}{8}\left(\frac{s\cos\theta}{M_{g}^{2}}\right)^{2}\Theta(\eta-\eta^{\prime})\Big[4(\eta^{\prime}-\eta_{0})\big(\left|x-x^{\prime}+\eta-\eta^{\prime}\right|+\left|x-x^{\prime}-\eta+\eta^{\prime}\right|\big)+\\ &(\eta-\eta^{\prime}+x-x^{\prime})|\eta-\eta^{\prime}+x-x^{\prime}|-(\eta^{\prime}-\eta+x-x^{\prime})|\eta^{\prime}-\eta+x-x^{\prime}|-\\ &(\eta-2\eta_{0}+\eta^{\prime}+x-x^{\prime})|\eta-2\eta_{0}+\eta^{\prime}+x-x^{\prime}|-(\eta-2\eta_{0}+\eta^{\prime}+x^{\prime}-x)|\eta-2\eta_{0}+\eta^{\prime}+x^{\prime}-x|\Big]+(\eta\mathrel{{\mathop{\longleftrightarrow}\limits^{\text{sym}}}}\eta^{\prime})~.\end{split}$ (B.1)
@@
$\begin{split}&\text{R}_{uu}\big(\eta,x;\eta^{\prime},x^{\prime}\big)=\frac{\langle u(\eta,x)u(\eta^{\prime},x^{\prime})\rangle-\langle u(\eta,x)\rangle\langle u(\eta^{\prime},x^{\prime})\rangle}{\sqrt{{\rm Var}\big[u(\eta,x)\big]{\rm Var}\big[u(\eta^{\prime},x^{\prime})\big]}}=\\ &-\frac{1}{8(\eta-\eta_{0})(\eta^{\prime}-\eta_{0})}\Big[4(\eta^{\prime}-\eta_{0})\big(\left|x-x^{\prime}+\eta-\eta^{\prime}\right|+\left|x-x^{\prime}-\eta+\eta^{\prime}\right|\big)+(\eta-\eta^{\prime}+x-x^{\prime})|\eta-\eta^{\prime}+x-x^{\prime}|\\ &-(\eta^{\prime}-\eta+x-x^{\prime})|\eta^{\prime}-\eta+x-x^{\prime}|-(\eta-2\eta_{0}+\eta^{\prime}+x-x^{\prime})|\eta-2\eta_{0}+\eta^{\prime}+x-x^{\prime}|\\ &-(\eta-2\eta_{0}+\eta^{\prime}+x^{\prime}-x)|\eta-2\eta_{0}+\eta^{\prime}+x^{\prime}-x|\Big]\Theta(\eta-\eta^{\prime})+(\eta\mathrel{{\mathop{\longleftrightarrow}\limits^{\text{sym}}}}\eta^{\prime})~.\end{split}$ (B.3)
@@ B.2 Autocorrelation of $v$
$\begin{split}&\langle v(\eta,x)v(\eta^{\prime},x^{\prime})\rangle-\langle v(\eta,x)\rangle\langle v(\eta^{\prime},x^{\prime})\rangle=\\ &\left(\frac{2s\sin\theta}{M_{g}^{2}}\right)^{2}\int\limits_{\eta_{0}}^{\eta}\mbox{d}\bar{\eta}\int\limits_{\eta_{0}}^{\eta^{\prime}}\mbox{d}\bar{\bar{\eta}}\int\limits_{-\infty}^{+\infty}\frac{\mbox{d}k_{x}}{2\pi}\,\frac{\sin(\omega_{k}(\eta-\bar{\eta}))}{\omega_{k}}\frac{\sin(\omega_{k}(\eta^{\prime}-\bar{\bar{\eta}}))}{\omega_{k}}e^{ik_{x}(x-x^{\prime})}\delta(\bar{\eta}-\bar{\bar{\eta}})=\\ &\left(\frac{2s\sin\theta}{M_{g}^{2}}\right)^{2}\Theta(\eta-\eta^{\prime})\int\limits_{\eta_{0}}^{\eta^{\prime}}\mbox{d}\bar{\eta}\int\limits_{-\infty}^{+\infty}\frac{\mbox{d}k_{x}}{2\pi}\,\frac{\sin(\omega_{k}(\eta-\bar{\eta}))\sin(\omega_{k}(\eta^{\prime}-\bar{\eta}))}{\omega_{k}^{2}}e^{ik_{x}(x-x^{\prime})}+(\eta\mathrel{{\mathop{\longleftrightarrow}\limits^{\text{sym}}}}\eta^{\prime})=\\ &\left(\frac{2s\sin\theta}{M_{g}^{2}}\right)^{2}\Theta(\eta-\eta^{\prime})\int\limits_{-\infty}^{+\infty}\text{d}k_{x}\;e^{ik_{x}(x-x^{\prime})}\Bigg(\frac{(\eta^{\prime}-\eta_{0})\cos{\left(\sqrt{k_{x}^{2}+m_{\rm FP}^{2}}(\eta-\eta^{\prime})\right)}}{4\pi(k_{x}^{2}+m_{\rm FP}^{2})}\\ &+\frac{\sin{\left(\sqrt{k_{x}^{2}+m_{\rm FP}^{2}}(\eta-\eta^{\prime})\right)}-\sin{\left(\sqrt{k_{x}^{2}+m_{\rm FP}^{2}}(\eta+\eta^{\prime}-2\eta_{0})\right)}}{8\pi(k_{x}^{2}+m_{\rm FP}^{2})^{3/2}}\Bigg)+(\eta\mathrel{{\mathop{\longleftrightarrow}\limits^{\text{sym}}}}\eta^{\prime})~.\end{split}$ (B.4)
@@
$\begin{split}&\int\limits_{-m_{\rm FP}}^{m_{\rm FP}}\text{d}u\,\dfrac{e^{-u\,q}}{m_{\rm FP}^{2}-u^{2}}\sin{\left(\sqrt{m_{\rm FP}^{2}-u^{2}}|p|\right)}\approx\\ &\quad\frac{1}{m_{\rm FP}}\Bigg[\pi{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sinh}(m_{\rm FP}|q|)-\sqrt{\frac{2\pi}{m_{\rm FP}|p|}}\left(1-\frac{q^{2}}{p^{2}}\right)^{1/4}\cos\Big(m_{\rm FP}\sqrt{p^{2}-q^{2}}+\frac{\pi}{4}\Big)\Bigg]~,\;\;\mbox{for $|p|>|q|$\,,\; $m_{\rm FP}|p|\gg 1$}~.\end{split}$ (B.11)
@@
    $\begin{split}&\langle v(\eta,x)v(\eta^{\prime},x^{\prime})\rangle-\langle v(\eta,x)\rangle\langle v(\eta^{\prime},x^{\prime})\rangle\\ &\approx\left(\frac{2s\sin\theta}{M_{g}^{2}}\right)^{2}\frac{(\eta^{\prime}-\eta_{0})\left((\eta-\eta^{\prime})^{2}-\Delta x^{2}\right)^{1/4}\cos{\left(\frac{\pi}{4}+m_{\rm FP}\sqrt{(\eta-\eta^{\prime})^{2}-\Delta x^{2}}\right)}}{2\sqrt{2\pi\,m_{\rm FP}^{3}}(\eta-\eta^{\prime})}\Theta(\eta-\eta^{\prime})\\ &\quad+\,(\eta\mathrel{{\mathop{\longleftrightarrow}\limits^{\text{sym}}}}\eta^{\prime})\end{split}$ (B.15)
@@
    $\begin{split}&\langle v(\eta,x)v(\eta^{\prime},x^{\prime})\rangle-\langle v(\eta,x)\rangle\langle v(\eta^{\prime},x^{\prime})\rangle\\ &\approx-\left(\frac{2s\sin\theta}{M_{g}^{2}}\right)^{2}\frac{\left((\eta+\eta^{\prime}-2\eta_{0})^{2}-\Delta x^{2}\right)^{3/4}\sin{\left(\frac{\pi}{4}+m_{\rm FP}\sqrt{(\eta+\eta^{\prime}-2\eta_{0})^{2}-\Delta x^{2}}\right)}}{4\sqrt{2\pi m_{\rm FP}^{5}}(\eta+\eta^{\prime}-2\eta_{0})^{2}}\Theta(\eta-\eta^{\prime})\\ &\quad+\,(\eta\mathrel{{\mathop{\longleftrightarrow}\limits^{\text{sym}}}}\eta^{\prime})~.\end{split}$ (B.16)
@@ B.3 Correlation between $u$ and $v$
$\begin{split}&\langle u(\eta,x)v(\eta^{\prime},x^{\prime})\rangle-\langle u(\eta,x)\rangle\langle v(\eta^{\prime},x^{\prime})\rangle=\\ &-\frac{2s^{2}\sin(2\theta)}{M_{g}^{4}}\Theta(\eta-\eta^{\prime})\int\limits_{\eta_{0}}^{\eta^{\prime}}\mbox{d}\bar{\eta}\int\limits_{-\infty}^{+\infty}\frac{\mbox{d}k_{x}}{2\pi}\,\frac{\sin(k_{x}(\eta-\bar{\eta}))\sin(\omega_{k}(\eta^{\prime}-\bar{\eta}))}{k_{x}\,\omega_{k}}e^{ik_{x}(x-x^{\prime})}+(\eta\mathrel{{\mathop{\longleftrightarrow}\limits^{\text{sym}}}}\eta^{\prime})\\ &=-\frac{2s^{2}\sin(2\theta)}{M_{g}^{4}}\Theta(\eta-\eta^{\prime})\int\limits_{-\infty}^{+\infty}\frac{\mbox{d}k_{x}}{2\pi m_{\rm FP}^{2}}\Bigg(\frac{\sin{\left(k_{x}(\eta-\eta^{\prime})\right)-\cos{\left(\sqrt{k_{x}^{2}+m_{\rm FP}^{2}}(\eta^{\prime}-\eta_{0})\right)}\sin{\left(k_{x}(\eta-\eta_{0})\right)}}}{k}\\ &+\frac{\cos{\left(k_{x}(\eta-\eta_{0})\right)}\sin{\left(\sqrt{k_{x}^{2}+m_{\rm FP}^{2}}(\eta^{\prime}-\eta_{0})\right)}}{\sqrt{k_{x}^{2}+m_{\rm FP}^{2}}}\Bigg)e^{ik_{x}(x-x^{\prime})}+(\eta\mathrel{{\mathop{\longleftrightarrow}\limits^{\text{sym}}}}\eta^{\prime})~,\end{split}$ (B.19)
@@
where sgn is the sign function with $\text{sgn(0)}=0$, $\Delta\eta^{\prime}=\eta^{\prime}-\eta_{0}>0~,\;\nu=x-x^{\prime}+\eta-\eta_{0}~,\;\tilde{\nu}=x-x^{\prime}-\eta+\eta_{0}$, with the auxiliary function
@@
    $\begin{split}&\langle u(\eta,x)v(\eta^{\prime},x^{\prime})\rangle-\langle u(\eta,x)\rangle\langle v(\eta^{\prime},x^{\prime})\rangle\\ &=-\frac{2s^{2}\sin(2\theta)}{M_{g}^{4}}\frac{\text{sgn}(\Delta x+(\eta-\eta^{\prime}))-\text{sgn}(\Delta x-(\eta-\eta^{\prime}))-2\cos{\left(m_{\rm FP}(\eta^{\prime}-\eta_{0})\right)}}{4m_{\rm FP}^{2}}\Theta(\eta-\eta^{\prime})\\ &\quad+\,(\eta\mathrel{{\mathop{\longleftrightarrow}\limits^{\text{sym}}}}\eta^{\prime})~.\end{split}$ (B.24)
@@
    $\begin{split}&\langle u(\eta,x)v(\eta^{\prime},x^{\prime})\rangle-\langle u(\eta,x)\rangle\langle v(\eta^{\prime},x^{\prime})\rangle\\ &\approx-\frac{2s^{2}\sin(2\theta)}{M_{g}^{4}}\frac{(\eta+\eta^{\prime}-2\eta_{0}-|\Delta x|)\sin{\left(\frac{\pi}{4}+m_{\rm FP}\sqrt{(\eta^{\prime}-\eta_{0})^{2}-(\eta-\eta_{0}-|\Delta x|)^{2}}\right)}}{2\sqrt{2\pi\,m_{\rm FP}^{5}}(\eta-\eta_{0}-|\Delta x|)[(|\Delta x|-(\eta-\eta^{\prime}))(\eta+\eta^{\prime}-2\eta_{0}-|\Delta x|)]^{1/4}}\Theta(\eta-\eta^{\prime})\\ &\quad+\,(\eta\mathrel{{\mathop{\longleftrightarrow}\limits^{\text{sym}}}}\eta^{\prime})~.\end{split}$ (B.25)
@@
[…]
@@
Finally, to compute the correlation of $u$ and $v$ from (4.6), we will also need the variance of $u$ and $v$ obtained above, given by (B.2) and (B.18), respectively.